# Unleashing the Potential of Large Language Models as Prompt Optimizers: An Analogical Analysis with Gradient-based Model Optimizers

### v1

@@ D.2 Experiment
Appendix E Prompts Optimized by Different Methods
@@
SGDM
Based on your current facing direction and any changes in direction, will following these step-by-step instructions, with explicit reference to direction and orientation, lead you back to the starting point?
APE
Will following the given set of instructions result in returning to the starting point?
APO
Based on the provided instructions, including the starting direction and the changes in direction throughout the sequence, determine if the person will return to the starting point.
OPRO
Think systematically and consider each step to determine the correct answer.
PE2
Carefully analyze the given instructions step by step and determine if you will return to the starting point.
GPO
Analyze the given step-by-step instructions in detail and determine if they will guide you back to the starting point. Carefully evaluate each instruction and vividly imagine the movements to make your decision.
@@
SGDM
What specific aspects of storytelling style, narrative structure, plot structure, character development, and narrative techniques should be considered when finding a movie similar to films like Pulp Fiction, Forrest Gump, Dances with Wolves, and The Usual Suspects? Analyze these elements in each movie’s construction and compare them against the given list to determine similarity. Focus on shared storytelling elements rather than thematic or genre similarities. Keep in mind that the most important factor in determining similarity might not be the overall theme or genre. Provide a single letter as your answer.
APE
Which of the following movies is most similar to the given movies based on their genre, themes, and plot elements? Choose the best option from the following choices:
APO
Which movie from the following options is most similar to the given movies, taking into account specific themes, plot, genre, characters, setting, tone, and style? Please provide a ranked list of the relevant elements mentioned above in determining similarity between movies. Consider the highest ranked element as the primary criteria and subsequent elements as secondary criteria in determining similarity. Additionally, consider the overall popularity and critical acclaim of the movies when making your selection. This will help ensure more accurate predictions.
OPRO
Which of the given movies is most similar to the listed options? Consider the movies Batman, The Usual Suspects, The Silence of the Lambs, and Jurassic Park. Choose the option that closely matches the given movies.
PE2
Select the movie option that is most closely related to the given list of movies after conducting a meticulous analysis.
GPO
Thoroughly analyze the themes, genres, and narrative elements of each film to identify the movie that best aligns with the provided options. Make a well-informed decision based on your comprehensive evaluation.
@@
SGDM
Count the number of items
APE
Develop a systematic approach to accurately determine the total number of items by counting each individual item separately and recording their corresponding quantities.
APO
What is the total count of mentioned items, considering each item individually without categorization and counting duplicates as separate items? Please provide the correct count as your output.
OPRO
How can we solve the problem by breaking it down step by step?
PE2
By employing a systematic and thorough approach, meticulously analyze each item in a step-by-step manner to precisely determine the total count.
GPO
Let’s break down the problem systematically by deconstructing it into individual steps and accurately computing the total number of objects.
@@
SGDM
Sort the given words, alphabetically, in case-sensitive order, considering the entire word, including all characters. The sorting should be done in ascending order based on the lowercase versions of the words, while preserving the original case and considering the entire word.
APE
How would you sort the given list of words in alphabetical order, considering both uppercase and lowercase letters?
APO
Sort the given list of words in lowercase alphabetical order, taking into account all characters including special characters and numbers. Ensure that the sorting process is case-insensitive and considers all characters, including special characters and numbers. Convert all letters to lowercase before sorting. Consider special characters and numbers in the sorting process.
OPRO
Arrange the given words in alphabetical order by considering only the first letter of each word and exclude punctuation or special characters.
PE2
Analyze the given words and provide a sorted list in alphabetical order.
GPO
Arrange the following words in alphabetical order:

### v2

@@ E.2 Experiment
Appendix F Prompts Optimized by Different Methods
@@
SGDM Based on your current facing direction and any changes in direction, will following these step-by-step instructions, with explicit reference to direction and orientation, lead you back to the starting point?
APE Will following the given set of instructions result in returning to the starting point?
APO Based on the provided instructions, including the starting direction and the changes in direction throughout the sequence, determine if the person will return to the starting point.
OPRO Think systematically and consider each step to determine the correct answer.
PE2 Carefully analyze the given instructions step by step and determine if you will return to the starting point.
GPO Analyze the given step-by-step instructions in detail and determine if they will guide you back to the starting point. Carefully evaluate each instruction and vividly imagine the movements to make your decision.
@@
SGDM What specific aspects of storytelling style, narrative structure, plot structure, character development, and narrative techniques should be considered when finding a movie similar to films like Pulp Fiction, Forrest Gump, Dances with Wolves, and The Usual Suspects? Analyze these elements in each movie’s construction and compare them against the given list to determine similarity. Focus on shared storytelling elements rather than thematic or genre similarities. Keep in mind that the most important factor in determining similarity might not be the overall theme or genre. Provide a single letter as your answer.
APE Which of the following movies is most similar to the given movies based on their genre, themes, and plot elements? Choose the best option from the following choices:
APO Which movie from the following options is most similar to the given movies, taking into account specific themes, plot, genre, characters, setting, tone, and style? Please provide a ranked list of the relevant elements mentioned above in determining similarity between movies. Consider the highest ranked element as the primary criteria and subsequent elements as secondary criteria in determining similarity. Additionally, consider the overall popularity and critical acclaim of the movies when making your selection. This will help ensure more accurate predictions.
OPRO Which of the given movies is most similar to the listed options? Consider the movies Batman, The Usual Suspects, The Silence of the Lambs, and Jurassic Park. Choose the option that closely matches the given movies.
PE2 Select the movie option that is most closely related to the given list of movies after conducting a meticulous analysis.
GPO Thoroughly analyze the themes, genres, and narrative elements of each film to identify the movie that best aligns with the provided options. Make a well-informed decision based on your comprehensive evaluation.
@@
SGDM Count the number of items
APE Develop a systematic approach to accurately determine the total number of items by counting each individual item separately and recording their corresponding quantities.
APO What is the total count of mentioned items, considering each item individually without categorization and counting duplicates as separate items? Please provide the correct count as your output.
OPRO How can we solve the problem by breaking it down step by step?
PE2 By employing a systematic and thorough approach, meticulously analyze each item in a step-by-step manner to precisely determine the total count.
GPO Let’s break down the problem systematically by deconstructing it into individual steps and accurately computing the total number of objects.
@@
SGDM Sort the given words, alphabetically, in case-sensitive order, considering the entire word, including all characters. The sorting should be done in ascending order based on the lowercase versions of the words, while preserving the original case and considering the entire word.
APE How would you sort the given list of words in alphabetical order, considering both uppercase and lowercase letters?
APO Sort the given list of words in lowercase alphabetical order, taking into account all characters including special characters and numbers. Ensure that the sorting process is case-insensitive and considers all characters, including special characters and numbers. Convert all letters to lowercase before sorting. Consider special characters and numbers in the sorting process.
OPRO Arrange the given words in alphabetical order by considering only the first letter of each word and exclude punctuation or special characters.
PE2 Analyze the given words and provide a sorted list in alphabetical order.
GPO Arrange the following words in alphabetical order: